OpenStudy (anonymous):

Will fan and medal! Use elimination to solve the story problem. Cristiano bought 24 baseball cards for $50. One type cost $1 per card, and the other cost $3 per card. Define the variables and write equations to find the number of each type of card he bought. Solve by using elimination.

OpenStudy (danjs):

Set up two equations total # cards = Total Cost =

OpenStudy (danjs):

x = 1 dollar cards y = 3 dollar cards total cards = 1 dollar cards + 3 dollar cards total cost = (1 dollar cards)*1 dollar + (3 dollar cards)* 3 dollar

OpenStudy (anonymous):

x+3y=50 5+3y=50 it means that 3y=45, y=15 next we see in first equation also 3y that are 45 , now we find x that is 5. it looks like ; 5+3(15)=50

OpenStudy (anonymous):

@TheSmartOne please help me!

OpenStudy (anonymous):

you understand me?